A force of 1.00x10² N is applied to a box with a mass of 2.00x10¹ kg, moving it across a floor. If the friction between the box and the floor is 2.00x10¹ N, what is the acceleration of the box?
1
Expert's answer
2020-11-25T10:45:46-0500

Let's apply the Newton's Second Law of Motion:


"\sum F_x=ma_x,""F_{appl}-F_{fr}=ma,""a=\dfrac{F_{appl}-F_{fr}}{m},""a=\dfrac{1.0\cdot10^2\ N-2.0\cdot10^1\ N}{2.0\cdot10^1\ kg}=4.0\ \dfrac{m}{s^2}."

Answer:

"a=4.0\ \dfrac{m}{s^2}."